# Inverse Hamiltonian reduction in type A and generalized slices in the affine Grassmannian

### v1

@@ 8.1. Inverse Hamiltonian reduction for finite W-algebras
Note $H^{2}_{G\times\mathbb{G}_{\alpha}}(\mathrm{Gr}^{0}_{-\alpha}\times S_{G,\mu+\alpha})=0$, as $G\times\mathbb{G}_{\alpha}$ acts freely on $\mathrm{Gr}^{0}_{-\alpha}\times S_{G,\mu+\alpha}$ with quotient $\overset{\circ}{\mathfrak{G}}_{\alpha}\times S_{\mu}$, the latter of which is homotopy equivalent to $S^{1}$, so that there is a unique quantization up to strongly $G\times\mathbb{G}_{\alpha}$-equivariant isomorphism.
@@
The reduction ${\mathcal{A}}_{G,\mu}/\!\!/\!\!/_{\chi_{\alpha}}\mathbb{G}_{\alpha}$ admits a quantum inverse Hamiltonian reduction given by $(\overset{\circ}{\mathfrak{G}}_{\alpha},{\bf m},{\mathcal{D}}^{\rm loc}_{\hbar,\alpha},\psi_{\mu,\alpha})$.
@@
By Corollary 8.4 we see that $((\overset{\circ}{\mathfrak{G}}_{\alpha},{\bf m})$ is inverse Hamiltonian reduction data for the equivariant Slodowy slices and by Proposition 8.8, $({\mathcal{D}}^{\rm loc}_{\hbar,\alpha},\psi_{\mu,\alpha})$ completes this to a quantum inverse Hamiltonian reduction. ∎
@@
(8.1.9) $A_{H\times\mathbb{G}_{\hbar}-{\rm fin}}\coloneqq\{a\in A\,|\,a\text{ is in a finite dimensional }H\times\mathbb{G}_{\hbar}\text{ representation }\}~{},$
@@
(8.1.11) ${\mathcal{W}}_{\mu}\cong{\mathcal{W}}_{G,\mu}^{G}~{}.$
@@ 8.2. Inverse Hamiltonian reduction for Harish Chandra bimodules
Consider the morphism, $M/\!\!/\!\!/_{\chi_{\mu}}N_{\chi_{\mu}}\rightarrow\big{(}M\otimes D_{\hbar(G)}\big{)}/\!\!/\!\!/_{(\chi_{\mu},0)}(N_{\chi_{\mu}}\times G)$, that sends a class $[m]\in(M/M_{\chi_{\mu}})^{N_{\chi}}$ to $[m\otimes 1]\in((M\otimes D_{\hbar}(G))/U_{\hbar}(\mathfrak{g})(M\otimes D_{\hbar}(G))_{\chi_{\mu}})^{N_{\chi_{\mu}}\times G}$.
@@
Suppose now that $A$ is an associative algebra object in $\rm HC_{\hbar}$. By unraveling definitions, we note that this is essentially the same as a $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$-algebra, $A$, flat over $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ and complete in the $\hbar$-adic topology, and equipped with a comoment map, i.e., a map of associative $\Phi^{\#}:U_{\hbar}(\mathfrak{g})\rightarrow A$ and an action of $G$ by algebra automorphisms such that the differential of the $G$ action agrees with the adjoint action from the comoment map. The preceding theorem also applies in this setting:

### v2

@@ 8.3. Inverse Hamiltonian reduction for finite W-algebras
Note $H^{2}_{G\times\mathbb{G}_{\alpha}}(\mathrm{Gr}^{0}_{-\alpha}\times S_{G,\mu+\alpha})=0$, as $G\times\mathbb{G}_{\alpha}$ acts freely on $\mathrm{Gr}^{0}_{-\alpha}\times S_{G,\mu+\alpha}$ with quotient $\overset{\circ}{\mathfrak{G}^{*}}_{\alpha}\times S_{\mu}$, the latter of which is homotopy equivalent to $S^{1}$, so that there is a unique quantization up to strongly $G\times\mathbb{G}_{\alpha}$-equivariant isomorphism.
@@
The reduction ${\mathcal{A}}_{G,\mu}/\!\!/\!\!/_{\chi_{\alpha}}\mathbb{G}_{\alpha}$ admits a quantum inverse Hamiltonian reduction given by $(\overset{\circ}{\mathfrak{G}^{*}}_{\alpha},{\bf m},{\mathcal{D}}^{\rm loc}_{\hbar,\alpha},\psi_{\mu,\alpha})$.
@@
By Corollary 8.4 we see that $((\overset{\circ}{\mathfrak{G}^{*}}_{\alpha},{\bf m})$ is inverse Hamiltonian reduction data for the equivariant Slodowy slices and by Proposition 8.10, $({\mathcal{D}}^{\rm loc}_{\hbar,\alpha},\psi_{\mu,\alpha})$ completes this to a quantum inverse Hamiltonian reduction. ∎
@@
[…]
@@
(8.3.5) ${\mathcal{W}}_{\mu}\cong({\mathcal{W}}_{G,\mu})^{G}~.$
@@ 8.4. Inverse Hamiltonian reduction for Harish Chandra bimodules
Consider the morphism, $M/\!\!/\!\!/_{\chi_{\mu}}N_{\chi_{\mu}}\rightarrow\big{(}M\otimes D_{\hbar(G)}\big{)}/\!\!/\!\!/_{(\chi_{\mu},0)}(N_{\chi_{\mu}}\times G)$, that sends a class $m]\in(M/M_{\chi_{\mu}})^{N_{\chi}}$ to $m\otimes 1]\in((M\otimes D_{\hbar}(G))/U_{\hbar}(\mathfrak{g})(M\otimes D_{\hbar}(G))_{\chi_{\mu}})^{N_{\chi_{\mu}}\times G}$.
@@
Suppose now that $A$ is an associative algebra object in $\rm HC_{\hbar}$. By unraveling definitions, we note that this is essentially the same as an associative $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$-algebra, $A$, flat over $\mathbb{\mathbb{C}}{{[\![\hbar]\!]}}$ and complete in the $\hbar$-adic topology, and equipped with a comoment map, i.e., a map of associative algebras $\Phi^{\#}:U_{\hbar}(\mathfrak{g})\rightarrow A$ and an action of $G$ by algebra automorphisms such that the differential of the $G$ action agrees with the adjoint action from the comoment map. The preceding theorem also applies in this setting: